# FactCheXcker: Mitigating Measurement Hallucinations in Chest X-ray Report Generation Models

### v1

@@ 4.4 Evaluation Metrics
Presence Hallucination. To evaluate hallucinations related to the identification of findings or devices, we employ Precision, Recall, F1-Score and Balanced Accuracy. These metrics measure the model’s ability to identify the ETT presence accurately without introducing false mentions. Since the FactCheXcker pipeline does not address false positive rate, we focus on the precision score. All the result metrics are included in the Appendix.
@@
Placement Hallucination. In the current setting, we use a rule-based approach to interpret ETT placement as “correct” or “incorrect” based on ETT measurement. In cases where a model-generated report identifies the presence of an ETT but lacks details of its placement or measurement, we assume by default that the placement is correct. We base this on the assumption that the report should explicitly mention any notable finding that needs attending. Then, we compute Precision, Recall, F1-Score, and Balanced Accuracy. The Appendix includes all the result metrics.
@@ 5.1 Analysis of Measurement Hallucinations
Figure 5: FactCheXcker Case Studies. Given an input image and an initial model-generated report, FactCheXcker analyzes the report and generates a query for the code generator, which synthesizes a program. This program is then executed using a Python interpreter to produce the output. The update module leverages this output to generate a final, updated report with accurate measurements.
@@ 5.2 FactCheXcker for ETT Presence
We evaluate FactCheXcker’s capability in improving ETT presence detection, which is a binary classification task determining whether an endotracheal tube is present in the chest X-ray image. Notably, we consider all false positives as object hallucinations, regardless if the model report contains an accompanied measurement. Table 4 presents a comparison between the original and updated models on the Precision score. The results demonstrate that FactCheXcker consistently enhances the ETT detection performance across all models. Model LLM-CXR shows the most significant improvement in precision with an increase of 625%. MedVersa and CheXagent achieved the highest precision scores after enhancement (0.80 and 0.76 respectively), demonstrating FactCheXcker’s ability to further improve models with strong baseline performance. These improvements indicate that FactCheXcker effectively enhances models’ ability to accurately identify the presence of endotracheal tubes while reducing false detections.
Note that in our current pipeline configuration, edits are made to the original report only if it mentions an endotracheal tube. Consequently, our analysis focuses on precision rather than recall, as the pipeline is not designed to address false negatives. A potential approach to mitigate this limitation involves running FactCheXcker on every image, independent of the report mentions. While this method may raise the false positive rate—dependent on the performance of the exists(“endotracheal tube”) quality. Further refinement of these modules could potentially enhance results by reducing false positives and improving overall accuracy.
@@ 5.3 FactCheXcker for ETT Measurement
To evaluate FactCheXcker’s effectiveness in mitigating measurement hallucinations, we analyze the measurement error on the instances where ETT is mentioned in both the original model-generated report and the ground truth report. As shown in Table 4, for all models, FactCheXcker achieves significant improvement in MAE score and the composite scoring that combines MAE and precision. FactCheXcker obtains an average improvement of 94% in the MAE score and an average improvement of 171% in the composite score. Specifically, models like VLCI, Cvt2distilgpt2, and LLM-CXR show remarkable improvements. The results demonstrate a significant reduction in measurement errors across all metrics, indicating FactCheXcker’s effectiveness in improving measurement accuracy for both correctly and incorrectly identified ETT cases.
@@
Figure 4 presents the failure rates across models, where a failure is defined as a measurement error exceeding 1.5 cm. Across all models, FactCheXcker demonstrates a clear improvement, significantly reducing the number of failed cases—nearly halving the failures on average. FactCheXcker achieved the most dramatic reduction with GPT4V, lowering the failure rate from 77.5% to 22.5%. Notably, FactCheXcker’s impact was particularly strong for models with high initial failure rates (above 70%), such as GPT4V, CheXagent, and LLM-CXR, reducing their failure rates to below 31%. This consistent reduction in error demonstrates FactCheXcker’s effectiveness in mitigating measurement hallucinations across diverse model architectures, establishing it as a valuable tool for enhancing the reliability of quantitative assessments in medical imaging models.
@@ 5.4 FactCheXcker for ETT Placement
We evaluate FactCheXcker’s effectiveness in assessing ETT placement correctness, which classifies whether an endotracheal tube is positioned correctly or requires adjustment due to being too high or too low. Table 4 compares the performance between the original and updated models and shows significant improvements in placement precision for all models with FactCheXcker, with an average improvement in placement precision from 0.67 to 0.77. Notably, for the LLM-CXR model, FactCheXcker increases placement precision from 0.62 to 0.79.
5.5 Case Studies
In Figure 5. we present examples illustrating FactCheXcker’s performance in different scenarios. FactCheXcker notably improves measurement accuracy. Specifically, in Case 2, it successfully detects the lower placement of the tube, which is positioned only 1.5 cm above the carina, demonstrating its effectiveness. Beyond its current capabilities, the FactCheXcker framework holds significant potential for expansion. A promising enhancement would be integrating a module for detecting and segmenting lesions, opacities, and nodules. Thus, the framework would be able to answer queries such as “measure the cardiac-thoracic ratio?”, or “get the diameter of the largest nodule in the posterior right lung base”. Moving forward, we will continually integrate additional tool modules into the framework to support an expanding range of diagnostic tasks.
@@ 6 Conclusion
Acknowledgment. Part of this effort was supported by Harvard Medical School – Seoul National University Hospital – Seoul National University College of Medicine Collaborative Research Program (HMS-SNUH-SNUCM Collaborative Research Award).
Appendix
@@ A Prompts
Generate Queries from Report Prompt You are an intelligent radiologist tasked with verifying key measurements in radiology reports. Task: Given a radiology report, your responsibility is to identify if it mentions any of [endotracheal tube] and formulate a measurement query to verify the results in the report. **Examples**: - Measure the distance between the carina and the endotracheal tube.” - Measure the diameter of the lung nodules in the upper left lung.” - Are there any lung masses present in the image?” **Output Requirements**: - Make sure the object is present in the image (positive examples). No pneumothorax and The patient has been extubated are two negative examples. - Only output measurement queries, and no validation queries. - Format each query as a numbered list with no new lines between the queries. - If there are no relevant queries, output an empty string ””.
@@
Generate Code from Queries Prompt You are an intelligent code generation agent responsible for transforming user queries into executable Python code using a specified API reference. **Task**: When given a query, generate Python code that addresses the query using the provided API. **Output Requirements**: - The generated code must be encapsulated in a function named get_result,’ which takes a parameter cxr_image of type CXRImage.’ - If the method find returns an empty list, the code must return the object that is not found like this: ”[object_name] not present in the image.” - After executing the code, return a string that answers the original query based on the results obtained. **Code Template**: Use the following structure for your code: “‘python def get_result(cxr_image: CXRImage): # Implement the logic to process the CXRImage and solve the query return result # Replace with the actual result based on the execution “‘ **API Reference**: {api_reference}
@@
API Reference class CXRObject: ”””A Python class containing an image object with bounding box information. Parameters ———- object_name : str The object name. bbox : Tuple[float, float, float, float] A tuple representing the bounding box in the format (left, lower, right, upper). The object is considered a point if left == right and lower == upper. Methods ——- is_point() -> bool Returns true if the objects bounding box is a point. get_center() -> Tuple[float, float] Returns the center of the objects bounding box or the point if its a single point. ””” class CXRImage: ”””A Python class containing an image related to a report as well as relevant information. Parameters ———- rid: str Id of report object reports: dict Reports of GroundTruth and models. image_path: str Full image path. original_size: List[int] The original WxH of the image. pixel_spacing: Tuple[float, float] Pixel spacing (mm) in the x- and y-directions. cache: ModuleCache An instance of ModuleCache to retrieve precomputed segmentation and measurement data. Methods ——- exists(object_name: str) -> bool Returns True if the object is found in the image, and False otherwise. find(object_name: str) -> List[CXRObject] Returns a list of CXRObjects found in the image matching the object_name. segment(object_name: str) -> CXRSegmentation Returns a segmentation map of the image based on the object specified. within(ob: CXRObject, region: CXRSegmentation) -> bool Returns true of if the object center is within the region. distance(obj_a: CXRObject obj_b: CXRObject) -> float Returns the distance (in cm) between the center of two objects in the image. diameter(obj: CXRObject) -> float Returns the diameter (in cm) of the CXRObject. dimensions(obj: CXRObject) -> Tuple[float, float] Returns the dimensions (in cm) of the CXRObject according to major axis x minor axis. width(segmentation: CXRSegmentation) -> float Returns the greatest width (in cm) of a segmentation. height(segmentation: CXRSegmentation) -> float Returns the greatest height (in cm) of a segmentation. filter(objects: [CXROBject], region: CXRSegmentation) -> List[CXRObject] Returns all objects with their center within the region. ””” class CXRSegmentation: ”””A Python class containing an anatomical region. Parameters ———- object_name : str The object name. segmentation_map : List[int, int] A binary segmentation map of the anatomical region. Methods ——- get_pixel_width() -> float Returns the widest pixel width of the segmentation get_pixel_height() -> float Returns the tallest pixel height of the segmentation ”””
@@
Update Report Prompt You are a highly skilled radiologist responsible for updating an existing radiology report based on a list of new results. **Task**: Using the provided list of new results, carefully revise the original report to reflect the new information. **Requirements**: - You must use all results in the list. - If any of the new results contradict details in the original report, update those details to align with the new results. - Retain the formatting, tone, and language of the original report as much as possible. - If the new results indicate that any sentence or phrase in the original report is incorrect, omit it from the updated version. - If the results contain no new measurements, you must keep the measurements in the original report. **Output Format**: Return the updated report using the following format: Updated report: ”{updated-report-here}”
@@
Extract ETT Mentions Prompt You are an experienced radiologist responsible for finding information on endotracheal tube (ET) placement in radiology reports. Your performance and accuracy are crucial for our patient care quality. To solve this task, perform the following steps: 1. Identify ET Tube Present: Determine if the report explicitly states that an ET tube is present. Note that mentions of ET tube removal or patient extubation indicate that the ET tube is no longer present. 2. Extract ET Tube Measurement: If an ET tube is present, extract its relative distance to the carina in centimeters (cm) if specified. Positive values indicate placement above the carina, while negative values indicate placement below the carina. 3. Determine ET Tube Placement: If an ET tube is present, determine if the report deems the placement correct or incorrect. If incorrect, categorize the placement as too low or too high,” if possible. If you cannot extract a specific category, use null”. There is no need to guess or invent a value. Adhere to the following rules: - Interpret measurements such as less than [x] cm as ”[x] cm”. - Interpret measurements such as ”[x]-[y] cm as the higher value, outputting ”[y] cm - Interpret measurements such as ”2. 0 cm as ”2.0 cm - Interpret terms like stable and unchanged as correct.” - If the report does not clarify whether the measurement is above or below the carina, assume it is above and provide a positive value. - If you do not find a specific measurement in centimeters (cm) or millimeters (mm), never infer or approximate a value, simply output null’. This is true even if the report specifies anatomical landmarks. First, write one sentence describing how you solved the task for each step. Finally, format your results as a JSON object using the following schema: “‘json { ET_present’: bool, ET_measurement’: float or null, ET_placement’: str or null, } “‘
@@
Extract Measurement Mentions Prompt You are an experienced radiologist responsible for finding information on object measurements in radiology reports. Your performance and accuracy are crucial for our quality of patient care. To solve this task, perform the following steps: 1. Identify measured object(s): list all objects that the reports include measurements for using concrete measurements in centimeters (cm) and/or millimeters (mm). Adhere to the following rules: - Do not include objects specified in qualitative descriptions or anatomical landmarks, such as incorrect position and projects into the stomach.” First, write one sentence describing how you solved the task for each step. Finally, format your answer as a comma-separated string, following this format: Reasoning: [1 sentence explaining your reasoning] Answer: object 1, object 2, etc., or an empty string if no objects are measured. **Examples:** Report: The tracheostomy tube ends 3.5 cm from the carina. There is a small apical right pneumothorax. Heart size is normal-the endotracheal tube projects into the T2 region. Reasoning: Only the tracheostomy tube contains a specific measurement in centimeters or millimeters. Answer: tracheostomy tube Report: Ill-defined nodule in the right upper lung measuring 1.3 x 1.4 cm. The endotracheal tube tip now measures approximately 4.6 cm above the carina-the tip of the right internal jugular vein catheter projects over the cavoatrial junction. Reasoning: Both the nodule and the endotracheal tube tip include specific measurements in centimeters or millimeters, which is why the answer consists of both objects but not the right internal jugular vein. Answer: nodule, endotracheal tube

### v2

@@ 4.4 Evaluation Metrics
Presence Hallucination. To evaluate hallucinations related to the identification of findings or devices, we employ Precision, Recall, F1-Score and Balanced Accuracy. These metrics measure the model’s ability to identify the ETT presence accurately without introducing false mentions. Since the FactCheXcker pipeline does not address false positive rate, we focus on the precision score. All the result metrics are included in the Supplementary Material.
@@
Placement Hallucination. In the current setting, we use a rule-based approach to interpret ETT placement as “correct” or “incorrect” based on ETT measurement. In cases where a model-generated report identifies an ETT but lacks details of its placement or measurement, we assume by default that the placement is deemed correct. We base this on the assumption that the report should explicitly mention any notable finding that needs attending. Then, we compute Precision, Recall, F1-Score, and Balanced Accuracy. The Supplementary Material includes all result metrics.
@@ 5.1 Analysis of Measurement Hallucinations
Figure 5: FactCheXcker Case Studies. (Left) Given an input image and an initial model-generated report, FactCheXcker generates queries for the code generator, which synthesizes a program. This program is then executed using a Python interpreter to produce the output. The update module leverages this output to generate an updated report with accurate measurements. (Right) Failure cases include tube misidentifications and obscured carina landmarks.
@@ 5.2 FactCheXcker for ETT Presence
We evaluate FactCheXcker’s capability in improving ETT presence detection, which is a binary classification task determining whether an endotracheal tube is present in the chest X-ray image. Notably, we consider all false positives as object hallucinations, regardless if the model report contains an accompanied measurement. Table 4 presents a comparison between the original and updated models on the Precision score. The results demonstrate that FactCheXcker consistently enhances the ETT detection performance across all models. Model LLM-CXR shows the most significant improvement in precision with an increase of 1024%. MedVersa and CheXagent achieved the highest precision scores after enhancement (0.80 and 0.76 respectively), demonstrating FactCheXcker’s ability to further improve models with strong baseline performance. These improvements indicate that FactCheXcker effectively enhances models’ ability to accurately identify the presence of endotracheal tubes while reducing false detections.
Note that in our current pipeline configuration, edits are made to the original report only if it mentions an endotracheal tube. Consequently, our analysis focuses on precision rather than recall, as the pipeline is not designed to address false negatives. A potential approach to mitigate this limitation involves running FactCheXcker on every image, independent of the report mentions. The results in the Supplementary Materials show that although this approach improves false negatives, the precision score significantly drops. The false positive rate depends on the “exists” modules. Further refinement of these modules could enhance results by reducing false positives and improving accuracy.
@@ 5.3 FactCheXcker for ETT Measurement
To evaluate FactCheXcker’s effectiveness in mitigating measurement hallucinations, we analyze the measurement error on the instances where ETT is mentioned in both the original model-generated report and the ground truth report. As shown in Table 4, for all models, FactCheXcker achieves significant improvement in MAE score and the composite scoring that combines MAE and precision. FactCheXcker obtains an average improvement of 135% in the MAE score and an average improvement of 186% in the composite score. Specifically, models like VLCI, Cvt2distilgpt2, and LLM-CXR show remarkable improvements. The results demonstrate a significant reduction in measurement errors across all metrics, indicating FactCheXcker’s effectiveness in improving measurement accuracy for both correctly and incorrectly identified ETT cases.
@@
Figure 4 presents the failure rates across models, where a failure is defined as a measurement error exceeding 1.5 cm. Across all models, FactCheXcker demonstrates a clear improvement, significantly reducing the number of failed cases—nearly halving the failures on average. FactCheXcker achieved the most dramatic reduction with GPT4V, lowering the failure rate from 77.5% to 22.5%. Notably, FactCheXcker’s impact was particularly strong for models with high initial failure rates (above 70%), such as GPT4V, CheXagent, and LLM-CXR, reducing their failure rates to below 20%. This consistent reduction in error demonstrates FactCheXcker’s effectiveness in mitigating measurement hallucinations across diverse model architectures, establishing it as a valuable tool for enhancing the reliability of quantitative assessments in medical imaging models.
@@ 5.4 FactCheXcker for ETT Placement
We evaluate FactCheXcker’s effectiveness in assessing ETT placement correctness, which classifies whether an endotracheal tube is positioned correctly or requires adjustment due to being too high or too low. Table 4 compares the performance between the original and updated models and shows significant improvements in placement precision for all models with FactCheXcker, with an average improvement in placement precision from 0.84 to 0.94. Notably, for the LLM-CXR model, FactCheXcker increases placement precision from 0.74 to 0.97.
5.5 Discussion
Case Studies. In Figure 5 we present examples illustrating FactCheXcker’s performance in a successful case and in four failure cases. Specifically, in the successful case, FactCheXcker detects and corrects the faulty ETT measurement from 6 cm to about 4 cm above the carina. Our radiologist team manually reviewed all failure cases (20 in total, 1.95% of all measurements). They identified three failure modes: tube misidentification (60%), obscured carina landmarks due to haziness (35%), and rotated images (5%). Training is needed for complex cases with multiple tubes.
Future Development. A promising enhancement would be the integration of a module to detect and segment lesions, opacities, and nodules. Thus, the framework would be able to answer queries such as “measure the cardiac-thoracic ratio”, or “get the diameter of the largest nodule in the posterior right lung base”. Moving forward, we will continually integrate additional tool modules into the framework to support an expanding range of diagnostic tasks.
@@ 6 Conclusion
Acknowledgment. Part of this effort was supported by the Harvard Medical School–Seoul National University Hospital–Seoul National University College of Medicine Collaborative Research Program (HMS-SNUH-SNUCM Collaborative Research Award). This award is part of the Boston-Korea Innovative Research Project, funded by the Korea Health Industry Development Institute (KHIDI) and the Ministry of Health and Welfare, Republic of Korea (Grant No. RS-2024-00403047).
Supplementary Material
@@ A Prompts
Generate Queries from Report Prompt You are an intelligent radiologist tasked with verifying key measurements in radiology reports. Task: Given a radiology report, your responsibility is to identify if it mentions any of [endotracheal tube] and formulate a measurement query to verify the results in the report. **Examples**: - "Measure the distance between the carina and the endotracheal tube." - "Measure the diameter of the lung nodules in the upper left lung." - "Are there any lung masses present in the image?" **Output Requirements**: - Make sure the object is present in the image (positive examples). "No pneumothorax" and "The patient has been extubated" are two negative examples. - Only output measurement queries, and no validation queries. - Format each query as a numbered list with no new lines between the queries. - If there are no relevant queries, output an empty string "".
@@
Generate Code from Queries Prompt You are an intelligent code generation agent responsible for transforming user queries into executable Python code using a specified API reference. **Task**: When given a query, generate Python code that addresses the query using the provided API. **Output Requirements**: - The generated code must be encapsulated in a function named get_result,’ which takes a parameter cxr_image of type CXRImage.’ - If the method find returns an empty list, the code must return the object that is not found like this: "[object_name] not present in the image." - After executing the code, return a string that answers the original query based on the results obtained. **Code Template**: Use the following structure for your code: “‘python def get_result(cxr_image: CXRImage): # Implement the logic to process the CXRImage and solve the query return result # Replace with the actual result based on the execution “‘ **API Reference**: {api_reference}
@@
API Reference class CXRObject: """A Python class containing an image object with bounding box information. Parameters ———- object_name : str The object name. bbox : Tuple[float, float, float, float] A tuple representing the bounding box in the format (left, lower, right, upper). The object is considered a point if left == right and lower == upper. Methods ——- is_point() -> bool Returns true if the objects bounding box is a point. get_center() -> Tuple[float, float] Returns the center of the objects bounding box or the point if its a single point. """ class CXRImage: """A Python class containing an image related to a report as well as relevant information. Parameters ———- rid: str Id of report object reports: dict Reports of "GroundTruth" and models. image_path: str Full image path. original_size: List[int] The original WxH of the image. pixel_spacing: Tuple[float, float] Pixel spacing (mm) in the x- and y-directions. cache: ModuleCache An instance of ModuleCache to retrieve precomputed segmentation and measurement data. Methods ——- exists(object_name: str) -> bool Returns True if the object is found in the image, and False otherwise. find(object_name: str) -> List[CXRObject] Returns a list of CXRObjects found in the image matching the object_name. segment(object_name: str) -> CXRSegmentation Returns a segmentation map of the image based on the object specified. within(ob: CXRObject, region: CXRSegmentation) -> bool Returns true of if the object center is within the region. distance(obj_a: CXRObject obj_b: CXRObject) -> float Returns the distance (in cm) between the center of two objects in the image. diameter(obj: CXRObject) -> float Returns the diameter (in cm) of the CXRObject. dimensions(obj: CXRObject) -> Tuple[float, float] Returns the dimensions (in cm) of the CXRObject according to major axis x minor axis. width(segmentation: CXRSegmentation) -> float Returns the greatest width (in cm) of a segmentation. height(segmentation: CXRSegmentation) -> float Returns the greatest height (in cm) of a segmentation. filter(objects: [CXROBject], region: CXRSegmentation) -> List[CXRObject] Returns all objects with their center within the region. """ class CXRSegmentation: """A Python class containing an anatomical region. Parameters ———- object_name : str The object name. segmentation_map : List[int, int] A binary segmentation map of the anatomical region. Methods ——- get_pixel_width() -> float Returns the widest pixel width of the segmentation get_pixel_height() -> float Returns the tallest pixel height of the segmentation """
@@
Update Report Prompt You are a highly skilled radiologist responsible for updating an existing radiology report based on a list of new results. **Task**: Using the provided list of new results, carefully revise the original report to reflect the new information. **Requirements**: - You must use all results in the list. - If any of the new results contradict details in the original report, update those details to align with the new results. - Retain the formatting, tone, and language of the original report as much as possible. - If the new results indicate that any sentence or phrase in the original report is incorrect, omit it from the updated version. - If the results contain no new measurements, you must keep the measurements in the original report. **Output Format**: Return the updated report using the following format: Updated report: "{updated-report-here}"
@@
Extract ETT Mentions Prompt You are an experienced radiologist responsible for finding information on endotracheal tube (ET) placement in radiology reports. Your performance and accuracy are crucial for our patient care quality. To solve this task, perform the following steps: 1. Identify ET Tube Present: Determine if the report explicitly states that an ET tube is present. Note that mentions of ET tube removal or patient extubation indicate that the ET tube is no longer present. 2. Extract ET Tube Measurement: If an ET tube is present, extract its relative distance to the carina in centimeters (cm) if specified. Positive values indicate placement above the carina, while negative values indicate placement below the carina. 3. Determine ET Tube Placement: If an ET tube is present, determine if the report deems the placement correct or incorrect. If incorrect, categorize the placement as "too low" or "too high," if possible. If you cannot extract a specific category, use "null". There is no need to guess or invent a value. Adhere to the following rules: - Interpret measurements such as "less than [x] cm" as "[x] cm". - Interpret measurements such as "[x]-[y] cm" as the higher value, outputting "[y] cm" - Interpret measurements such as "2. 0 cm" as "2.0 cm" - Interpret terms like "stable" and "unchanged" as "correct." - If the report does not clarify whether the measurement is above or below the carina, assume it is above and provide a positive value. - If you do not find a specific measurement in centimeters (cm) or millimeters (mm), never infer or approximate a value, simply output null’. This is true even if the report specifies anatomical landmarks. First, write one sentence describing how you solved the task for each step. Finally, format your results as a JSON object using the following schema: “‘json { ET_present’: bool, ET_measurement’: float or null, ET_placement’: str or null, } “‘
@@
Extract Measurement Mentions Prompt You are an experienced radiologist responsible for finding information on object measurements in radiology reports. Your performance and accuracy are crucial for our quality of patient care. To solve this task, perform the following steps: 1. Identify measured object(s): list all objects that the reports include measurements for using concrete measurements in centimeters (cm) and/or millimeters (mm). Adhere to the following rules: - Do not include objects specified in qualitative descriptions or anatomical landmarks, such as "incorrect position" and "projects into the stomach." First, write one sentence describing how you solved the task for each step. Finally, format your answer as a comma-separated string, following this format: Reasoning: [1 sentence explaining your reasoning] Answer: object 1, object 2, etc., or an empty string if no objects are measured. **Examples:** Report: The tracheostomy tube ends 3.5 cm from the carina. There is a small apical right pneumothorax. Heart size is normal-the endotracheal tube projects into the T2 region. Reasoning: Only the tracheostomy tube contains a specific measurement in centimeters or millimeters. Answer: tracheostomy tube Report: Ill-defined nodule in the right upper lung measuring 1.3 x 1.4 cm. The endotracheal tube tip now measures approximately 4.6 cm above the carina-the tip of the right internal jugular vein catheter projects over the cavoatrial junction. Reasoning: Both the nodule and the endotracheal tube tip include specific measurements in centimeters or millimeters, which is why the answer consists of both objects but not the right internal jugular vein. Answer: nodule, endotracheal tube